OpenStudy (anonymous):

Help!!!!! Evaluate the indefinite integral 20tan^6x

hartnn (hartnn):

1) split up tan^6 as tan^4 . tan^2 2)write tan^2 = 1+sec^2 3) put t= tan x

hartnn (hartnn):

sorry, tan^2 = sec^2-1

hartnn (hartnn):

and 3) t= tan x only in the integral having sec^2x

OpenStudy (anonymous):

so \[20\int\limits_{}^{}\tan ^{4}x(\sec ^{2}x-1)\]

hartnn (hartnn):

do same thing for tan^4x dx

hartnn (hartnn):

yes, go on

OpenStudy (anonymous):

now i split tan^4x again?

hartnn (hartnn):

first do t^4(s^2-1) = t^4s^2 - t^4 so in 1st integral, t^4s^2 put tan x =t and split up t^4 as t^2.t^2

hartnn (hartnn):

separate the 2 integrals... did u understand that ? or confused because i wrote t and s for tan and sec ?

OpenStudy (anonymous):

no i get that

OpenStudy (anonymous):

so t^4s^2-t^2*t^2

hartnn (hartnn):

solve the 2 integrals separately...else you will get confused, because the 2nd integral is again going to split up into 2

OpenStudy (anonymous):

alright i see that. but where do i go from here. u substitution of the first part?

hartnn (hartnn):

yes

OpenStudy (anonymous):

so \[4\tan ^{5}-(20\int\limits_{}^{}\tan ^{4}x)\]

hartnn (hartnn):

yeah, solve t^4 in same manner.

OpenStudy (anonymous):

so \[\tan ^{5}x-\frac{ 20 }{ 3 }\tan ^{3}x+20tanx-20x\]

OpenStudy (anonymous):

+ C of course

hartnn (hartnn):

be careful with the signs..

OpenStudy (anonymous):

Those should be the correct signs according to the answers section in my book. Thanks for your help I was really struggling with this question

hartnn (hartnn):

oh, okk :) i though +20/3 will come, but anyways, welcome ^_^